# A Theory of LLM Sampling: Part Descriptive and Part Prescriptive

### v3

@@ Abstract
A Theory of LLM Sampling:
@@
Sarath Sivaprasad1, Pramod Kaushik2, Sahar Abdelnabi3, Mario Fritz1 1CISPA Helmholtz Center for Information Security  2TCS Research, Pune  3Microsoft {sarath.sivaprasad, fritz}@cispa.desaabdelnabi@microsoft.compramod.kaushik@tcs.com
@@ 1 Introduction
Decision making is a challenging task which often requires choosing an option from a vast set of possibilities (Mattar and Lengyel, 2022; Ross et al., 2023). In many real world cases deliberating on these innumerable options to decide on the action is computationally prohibitive, due to which, agents employ heuristics to sample their options Gigerenzer and Gaissmaier (2011). For instance, humans (and animals) are shown to deliberate only on a few options that are selected based on a heuristics guided by possibility (how statistically likely an option is) and utility (the value associated to the option) Bear et al. (2020); Mattar and Daw (2018). LLMs, though widely benchmarked as ‘system-1’ and characterised by their reliance on heuristics, have not been studied systematically for the heuristics that drive their sampling.
Figure 1: From left to right: when sampling on a concept, the LLM appears to account for the statistical likelihood ($A(C)$) and prescriptive norm ($I(C)$) of the concept. Consequently, the sampled distribution exhibits a shift (shown as $\alpha$) away from the true distribution in the direction of the ideal (right most plot).
We systematically study this sampling heuristics and show that they resemble that of human decision-making. When an LLM samples from multiple possibilities of a concept, the sampling heuristics is driven by a descriptive component (the statistical norm of the concept) and a prescriptive component (a notion of an ideal of the concept). These norms on the concept are learned in context or in pre-training. We define response sampling as the process by which the model probabilistically selects outputs from a distribution of potential options. A descriptive component represents what is statistically likely for a concept, reflecting the occurrence or probability of options. A prescriptive component is an implicit standard of what is considered ideal, desirable, or a valued option of a concept.
@@
Prescriptive component in options considered by humans is shown to be driven by their concept prototypicality which has a prescriptive component (e.g., a prototypical teacher is one that teaches well). We study prototypicality and sampling in humans and LLMs to make initial explorations into the cause of this deviation. Given that the training of LLMs is auto-regressive, it is not trivial that the heuristics of LLMs also behave as if they have a value component. In short, we make the following contributions:
@@
  • We illustrate the proposed theory using concept prototypicality. We show that the ideal notion in LLMs might not align with the value system of humans even though both LLMs and humans seem to share the same heuristic components.
@@ 2 Related Work
Earlier work that examined the mechanisms by which LLMs generate outputs considers that they produce coherent text by probabilistically assembling language patterns without ‘genuine understanding’ Bender et al. (2021). But, later investigations have demonstrated that LLMs can develop internal, structured representations of the environment Li et al. (2022). They even exhibit an understanding of semantic structures when trained on programming languages, indicating a capacity for meaningful text processing and generationJin and Rinard (2023). This has sparked interest within the community to explore the mechanisms governing output generation in LLMs through the lens of cognitive science and related disciplines.
Recent work indicates that LLM agents despite understanding the notion of probabilities struggle with probability sampling Gu et al. (2024). They do not fully represent the statistics, hindering their effectiveness in generating samples that align with expected probabilistic patterns. Our paper provides a systematic framework that explains the sampling behaviour of LLMs which can potentially explain the different biases shown by LLMs (more in Appendix C).
Understanding LLMs as ‘System-1’: Reasoning has been broadly characterized as a two-step process involving quick ‘System-1’ thinking and a more deliberate ‘System-2’ reasoning (Kahneman, 2011). Large Language Models (LLMs) have been conceptually likened to System-1 due to their heuristic-driven and non-deliberative output generation (Yao et al., 2023). In fact, recent studies show overlaps in errors made by LLMs and humans in System-1 reasoning tasks, indicating that both might rely on similar heuristics for rapid decision-making (Dasgupta et al., 2022). We study the convergence of heuristics between LLMs and humans and propose a theory for LLM sampling.
Previous research mainly uses sampling for tasks like action generation and decision making rather than to explicitly study the sampling mechanisms of LLMs (Hazra et al., 2023; Shah et al., 2023; Suri et al., 2023). Our work aims to fill this gap by investigating the heuristics driving LLMs’ response sampling, which could provide a deeper understanding of their decision-making processes.
@@ 3 Theory of LLM sampling
LLMs (without CoT Wei et al. (2022)) do not employ explicit deliberation for generating output. Hence, understanding the heuristics driving their sampling is key to explaining their performance. We examine the sampling mechanisms of LLMs in the light of human cognitive theory and propose a theory for LLM sampling:
@@ 3.2 Sampling in relation to existing concepts
Examples of this framework is shown in Figure 2. Consider the number of concepts for which sample falls on the ideal side of the average is $n$ and the total number of concepts evaluated is $n_{total}$. The Binomial test is used to determine if $n$ is significantly different from what would be expected by chance, assuming a null hypothesis where the probability $p$ of a sample being on the ideal side is 0.5. The $p$-value obtained from the binomial test is used to assess significance. $p<0.05$ shows presence of prescriptive norm across concepts.
@@
Comparing with human studies: The setting described in Sections 3.1 and  3.2 is inspired by similar evaluation in humans (Bear et al., 2020; Phillips et al., 2019; Bear and Knobe, 2017). We scale the experiments to show higher statistical significance. In appendix, we show replication of the exact setting of human studies (Table 4) on LLM (Table 5) to make a direct one to one comparison using the respective alphas (Figure 6 (left)). The conclusion follows that the heuristics of sampling converges in LLMs and humans but there is no guarantee that the prescriptive component aligns. This causes deviation of sample from the statistical likelihood in unforeseen directions, an interesting direction for future research in fairness and alignment.
@@ 4 Experiments and Results
In this section, we present two key experiments and a case study. First, we present a constrained setting to test the validity of the proposed theorem following the method in Section 3.1. Second, we evaluate the presence of prescriptive and descriptive components in sampling for concepts learned in training following Section 3.2. Our results show significant evidence for the proposed theory. We test on the instruction-tuned models of GPT-4 (Achiam et al., 2023), GPT-3.5-Turbo (Brown et al., 2020), Claude (Anthropic, 2024), Mixtral-8x7B (Jiang et al., 2024), Mistral-7B (Jiang et al., 2023), and both pretrained and instruction tuned models from the family of Llama-2 and 3 models (Touvron et al., 2023) B. Unless mentioned otherwise, we report results for GPT-4 in the main text and the results for other models in the Appendix. The complete text used in the prompts for each experiment is given in the Appendix I, L, Q and O respectively.
@@ 4.1 Sampling in relation to a novel concept
Results: Table 1 shows the result for the mean of the hundred runs for the uni-modal and bi-modal input distributions, each with three different $C_{v}$. Firstly, across the six settings, $A(C)$ approximately coincide with the true distribution average ($C_{\mu}=45$). For a neutral prescriptive norm (also for no prescriptive norm as shown later), $S(C)\approx A(C)\approx C_{\mu}$ and the input distribution and $S(C)$ do not differ significantly, $\textit{p}=0.52$. This shows that the sampling is driven solely by statistical considerations when no “ideal” notion is given.
@@
Figure 3: Variation of mean of $S(C)$ with changing prescriptive value.
@@
We observe statistically significant results for most other evaluated LLMs. Results for GPT-4 (with temperature set to zero), GPT-3.5-Turbo, Claude, Mixtral-8x7B, Mistral-7B, and Llama models are in Appendix N. For example, Claude-Opus, with a negative and positive $C_{v}$, $S(C)$ is statistically significant from $A(C)$ with p<.$<.001$.
uni-modal bi modal
@@ 4.2 Sampling in relation to existing concepts
In this experiment, the statistics $C_{\mu}$ and value system $C_{v}$ for a concept $C$ are implicit in the LLM and unknown to us. We empirically evaluate the proposed theory on 500 different concepts ($C$) spanning 10 different domains. The full list of concepts are in the Appendix Q. For each concept, we first ask the model to report its notion of (a) the average ($A(C)$) and (b) the ideal ($I(C)$) for a given concept $C$. We then give a sample prompt for concept $C$ to get (c) sample ($S(C)$). These prompts are given in independent contexts. To get these values, we use a prompt similar to the questions used in human studies (Bear et al., 2020). For example, to get the average, ideal, and the sample on the concept of ‘TV watching hours of people’, we use the following prompts:
@@
Table 2: Model Comparison across LLMs showing influence of the prescriptive component in existing concepts. The table shows a larger influence of prescriptive norms for larger model sizes and higher for RLHF compared to pretrained-only models.
@@
Case study for medical recovery time
: Deviation of a sample towards the prescriptive norm can help explain some biases of LLMs. To illustrate this, we present a case study based on a real-world scenario. The LLM agent is assigned the role of a doctor and asked to take a decision on the discharge time of a patient based on a list of symptoms. Here the action space is the positive rational numbers (number of weeks). Once the LLM gives a recovery time we also get self reported average and ideal recovery time from the LLM.
@@
The ideal value given by the LLM, is lower than the average value in 30 of the 35 symptoms. This implies that the sample is often pulled below the average. This finding indicates that LLMs’ decision-making regarding patient recovery times is compromised by a prescriptive component, which has significant implications for clinical decision-making, resource allocation in hospitals, and potential risks to patient safety. The full list of the symptoms and the exact prompts used is given in the Appendix M.
@@ 5 Prescriptive component in concept prototypes
One of the basic characteristics of System-1 is that it represents concepts with prototypical examples Kahneman (2011). In humans, though a prototype is often understood as the most typical/representative member of a concept (Murphy, 2004), they are found to embody both statistical regularities and goal-oriented ideals within a concept (Barsalou, 1985). For instance, a ‘Robin’ might be considered a prototype of the concept ‘Bird’, as it shares many common features with most birds with high occurrence (statistics), and has the ability to fly (a value expected of birds), making it a prototypical example of the concept ‘bird’ (Smith and Medin, 1981). For this reason, penguin-a flightless bird, is less prototypical bird than ‘Robin’. Prototypicality defines normality of a concept that drive the sampling (Appendix  F).
Unlike humans, it is not clear whether LLMs rely on concept prototypes for sampling. But since the sampling heuristics of LLMs converge with humans, it is interesting to investigate concept prototypicality in LLMs. We do not claim that LLMs output is prototype driven, but make an initial exploration in this direction using the exact setting as in Bear and Knobe (2017). That is, we use eight concepts, and for each concept $C$, six different exemplars. Exemplars are short descriptions of items of a concept. For instance, for the concept of ‘High-school teacher’, the first exemplar is as follows: ‘A 30-year-old woman who basically knows the material she is teaching but is relatively uninspiring, boring to listen to, and not particularly fond of her job.’
Similar to experiment protocol in Bear and Knobe (2017), LLMs rate each exemplars on three dimensions: average, ideal, and the prototypicality of the exemplar. Prototypicality score is derived by averaging three entities, which measure the degree to which the given prototype is a “good example”, “paradigmatic example”, or “prototypical example” Bear and Knobe (2017). The LLM is asked to rate on a 7-point scale ranging from not at all average/ideal/‘good example’, which has a score of 0, to completely average/ideal/‘good example’, with a score 7. The full set of concepts and exemplars are in Appendix O.
@@
Evaluating across different LLMs, we obtain the following results: Llama-3-7b (binomial $p$ = 0.003), Mixtral-8x7B (binomial $p$ = 0.05), GPT3.5-turbo (binomial $p$ < 0.001), Claude (binomial $p$ < 0.001), Mistral (binomial $p$ = 0.0019), indicating the effect of prescriptive norms in prototypes of concepts. The complete set of results for every exemplar is given in Appendix P. This experiment is an initial exploration, finding that LLMs’ concept of prototypes is influenced not only by statistical averages but also by an underlying prescriptive norm. These findings suggest that the LLM’s judgment of what constitutes a typical or prototypical example is systematically biased toward idealized representations calling for further investigations in this direction.

### v4

# A Theory of Response Sampling in LLMs: Part Descriptive and Part Prescriptive

@@ Abstract
A Theory of Response Sampling in LLMs:
@@
Sarath Sivaprasad1, Pramod Kaushik2, Sahar Abdelnabi3, Mario Fritz1 1CISPA Helmholtz Center for Information Security  2TCS Research, Pune  3Microsoft {sarath.sivaprasad, fritz}@cispa.depramod.kaushik@tcs.comsaabdelnabi@microsoft.com
@@ 1 Introduction
Decision making is a challenging task which often requires choosing an option from a vast set of possibilities (Mattar and Lengyel, 2022; Ross et al., 2023). In many real world cases deliberating on these innumerable options to decide on the action is computationally prohibitive, due to which, agents employ heuristics to sample their options Gigerenzer and Gaissmaier (2011). For instance, humans (and animals) are shown to deliberate only on a few options that are selected based on a heuristics guided by possibility (how statistically likely an option is) and utility (the value associated to the option) Bear et al. (2020); Mattar and Daw (2018). While LLMs are often described as ‘System-1’ (Appendix A) agents, characterised by their reliance on heuristics, the mechanism governing their response sampling remains under-explored.
Figure 1: From left to right: when sampling on a concept, the LLM appears to account for the statistical likelihood ($A(C)$) and prescriptive norm ($I(C)$) of the concept. Consequently, the sample distribution exhibits a shift (shown as $\alpha$) away from the true distribution in the direction of the ideal (right most plot).
We define response sampling as the process by which the LLM agent probabilistically selects outputs from a distribution of potential options (refer Appendix A for all formal definitions). We systematically study this sampling heuristics and show that they resemble that of human decision-making. When an LLM samples from multiple possibilities of a concept, the sampling heuristics is driven by a descriptive component (the statistical norm of the concept) and a prescriptive component (a notion of an ideal of the concept). A descriptive component represents what is statistically likely for a concept, reflecting the occurrence or probability of options. A prescriptive component is an implicit standard of what is considered ideal, desirable, or a valued option of a concept. These norms on the concepts can be learned both in-context or in pre-training.
@@
Human decision-making is often guided by concept prototypicality Murphy (2004), which incorporates both descriptive (statistically common) and prescriptive (value) components (Barsalou, 1985) - (e.g., while most teachers may exhibit certain average competence, the prototypical teacher is often imagined to teach well). We make initial investigation to show that LLMs’ prototypicality has these two components and hypothesis its connection to sampling. In short, we make the following contributions:
@@
  • We demonstrate that LLMs’ prototypical representations of concepts systematically incorporate prescriptive norms, showing initial evidence that their judgments of ’typical’ examples are biased toward idealized versions, similar to human notion of prototypicality.
@@ 2 Related Work
Earlier work that examined the mechanisms by which LLMs generate outputs suggests that they produce coherent text by probabilistically assembling language patterns without ‘genuine understanding’ Bender et al. (2021). But, later investigations have demonstrated that LLMs can develop internal, structured representations of the environment Li et al. (2023). They even exhibit an understanding of semantic structures when trained on programming languages, indicating a capacity for meaningful text processing and generation Jin and Rinard (2024). This has sparked interest within the community to explore the mechanisms governing output generation in LLMs through the lens of cognitive science and related disciplines.
Recent work indicates that LLM agents despite understanding the notion of probabilities struggle with probability sampling Gu et al. (2025). They do not fully represent the statistics, i.e. poor at generating samples that align with expected probabilistic patterns. Our paper provides a systematic framework that explains the components in samples of LLMs. This can potentially explain the different biases shown by LLMs (more in Appendix C).
Understanding LLMs as ‘System-1’: Reasoning has been broadly characterized as a two-step process involving quick ‘System-1’ thinking and a more deliberate ‘System-2’ reasoning (Kahneman, 2011). Large Language Models (LLMs) have been conceptually likened to System-1 due to their heuristic-driven and non-deliberative output generation (Yao et al., 2023). In fact, recent studies show overlaps in errors made by LLMs and humans in System-1 reasoning tasks, indicating that both might rely on similar heuristics for rapid decision-making (Lampinen et al., 2024). We study the convergence of heuristics between LLMs and humans and propose a theory for LLM sampling.
Previous research mainly uses sampling for tasks like action generation and decision making rather than to explicitly study the sampling mechanisms of LLMs (Hazra et al., 2024; Shah et al., 2023; Suri et al., 2024). Our work aims to fill this gap by investigating the heuristics driving LLMs’ response sampling, which could provide a deeper understanding of their decision-making processes.
@@ 3 Theory of LLM sampling
When faced with numerous possible actions, where deliberating on each option is computationally prohibitive, humans inherently resort to forming a finite consideration set of options using heuristics Phillips et al. (2019). Cognitive studies characterize this heuristic-based filtering as ‘System-1’ thinking: fast, automatic, and intuition-driven Kahneman (2011); Gigerenzer and Gaissmaier (2011). Such heuristics effectively reduce the cognitive load on deliberative processes (‘System-2’), by selecting a manageable subset of options for further deliberation. In humans, these heuristics are guided primarily by two factors: the statistical likelihood of options and their perceived value Bear et al. (2020).
In LLMs, reasoning mechanisms such as CoT Wei et al. (2022) and other explicit reasoning models like GPT-o3 OpenAI (2024) and Deepseek-r1 Guo et al. (2025) are likened to explicit deliberation (‘System-2’), while the default mechanism is likened to heuristics driven ‘System-1’ Li et al. (2025). Hence, understanding the heuristics driving their sampling is key to explaining their performance. We examine the sampling mechanisms of LLMs in the light of this human cognitive theory and propose a theory for LLM sampling:
@@ 3.2 Sampling in relation to existing concepts
Examples of this framework are shown in Figure 2. Consider the number of concepts for which sample falls on the ideal side of the average is $n$ and the total number of concepts evaluated is $n_{total}$. The Binomial test is used to determine if $n$ is significantly different from what would be expected by chance, assuming a null hypothesis where the probability $p$ of a sample being on the ideal side is 0.5. The $p$-value obtained from the binomial test is used to assess significance. $p<0.05$ shows presence of prescriptive norm across concepts.
@@
The deviation metric $\alpha$ is defined as a directional measure quantifying the shift from the statistical norm $A(C)$ toward the prescriptive norm $I(C)$. However, when the average and ideal values are equal (i.e., $A(C)=I(C)$), the directional term sign($A(C)-I(C)$) becomes zero, leading to an undefined $\alpha$ and $\hat{\alpha}$. These degenerate cases are excluded from the $\alpha$ based analysis, as no directional deviation can be computed. Such analysis is meaningful when there is a difference between statistical and prescriptive components.
Comparing with human studies: The setting described in Sections 3.1 and  3.2 is inspired by similar evaluation in humans (Bear et al., 2020; Phillips et al., 2019; Bear and Knobe, 2017). We scale the experiments to show higher statistical significance. In appendix, we show replication of the exact setting of human studies (Table 4) on LLM (Table 5) to make a direct one to one comparison using the respective $\alpha$s (Figure 6 (left)). The conclusion follows that the heuristics of sampling converges in LLMs and humans but the degree of deviation of sample towards the prescriptive norm does not align. This causes deviation of sample from the statistical likelihood in unforeseen degrees, an interesting direction for future research in fairness and alignment.
@@ 4 Experiments and Results
In this section, we present two key experiments and a case study. First, we present a constrained setting to test the validity of the proposed theorem following the method in Section 3.1. Second, we evaluate the presence of prescriptive and descriptive components in sampling for concepts learned in training following Section 3.2. Our results show significant evidence for the proposed theory. We test on the instruction-tuned models of GPT-4 (Achiam et al., 2023), GPT-3.5-Turbo (Brown et al., 2020), Claude (Anthropic, 2024), Mixtral-8x7B (Jiang et al., 2024), Mistral-7B (Jiang et al., 2023), and both pretrained and instruction tuned models from the family of Llama-2 and 3 models (Touvron et al., 2023) B. Unless mentioned otherwise, we report results for GPT-4 in the main text and the results for other models in the Appendix. The complete text used in the prompts for each experiment is given in the Appendix I, M, R and P respectively.
@@ 4.1 Sampling in relation to a novel concept
Results: Table 1 shows the result for the mean of the hundred runs for the uni-modal and bi-modal input distributions, each with three different $C_{v}$. Firstly, across the six settings, $A(C)$ approximately coincide with the true distribution average ($C_{\mu}=45$). For a neutral prescriptive norm (also for no prescriptive norm as shown later), $S(C)\approx A(C)\approx C_{\mu}$ and the input distribution and $S(C)$ do not differ significantly, $\textit{p}=0.52$. Given the non-significant difference ($\textit{p}=0.52$), the result is consistent with the hypothesis that sampling reflects the input distribution’s statistical properties when no “ideal” is specified.
@@
Figure 3: Variation of mean of $S(C)$ with changing prescriptive value. X-axis shows the different prescriptive values and Y-axis shows the sample value. Sample is directly proportional to prescriptive value. Here $C_{\mu}=45,$ approximately equal to $A_{C}$.
@@
uni-modal bi-modal
@@
We observe statistically significant results for most other evaluated LLMs. Results for GPT-4 (with temperature set to zero), GPT-3.5-Turbo, Claude, Mixtral-8x7B, Mistral-7B, and Llama models are in Appendix O. For example, Claude-Opus, with a negative and positive $C_{v}$, $S(C)$ is statistically significant from $A(C)$ with p<.$<.001$.
@@ 4.2 Sampling in relation to existing concepts
In this experiment, the statistics $C_{\mu}$ and value system $C_{v}$ for a concept $C$ are implicit in the LLM and unknown to us. We empirically evaluate the proposed theory on 500 different concepts ($C$) spanning 10 different domains. The full list of concepts are in the Appendix R. For each concept, we first ask the model to report its notion of (a) the average ($A(C)$) and (b) the ideal ($I(C)$) for a given concept $C$. We then give a sample prompt for concept $C$ to get (c) sample ($S(C)$). These prompts are given in independent contexts. To get these values, we use a prompt similar to the questions used in human studies (Bear et al., 2020). For example, to get the average, ideal, and the sample on the concept of ‘TV watching hours of people’, we use the following prompts:
@@
Table 2: Model comparison across LLMs showing influence of the prescriptive component in existing concepts.The fraction indicates the proportion of concepts within each domain for which the LLM’s sampled value deviates from the average in the direction of the ideal. The table shows a larger influence of prescriptive norms for larger model sizes and higher for RLHF compared to pretrained-only models.
@@
Case study for medical recovery time:
Deviation of a sample towards the prescriptive norm can help explain some biases of LLMs. To illustrate this, we present a case study based on a real-world scenario. The LLM agent is assigned the role of a doctor and asked to take a decision on the discharge time of a patient based on a list of symptoms. Here the action space is the positive rational numbers (number of weeks). Once the LLM gives a recovery time we also get self reported average and ideal recovery time from the LLM. The term self-reported average refers explicitly to the average values directly provided by the LLM itself when prompted to report average.
@@
The ideal value given by the LLM, is lower than the average value in 30 of the 35 symptoms. This implies that the sample is often pulled below the average. This finding indicates that LLMs’ decision-making regarding patient recovery times is compromised by a prescriptive component, which has significant implications for clinical decision-making, resource allocation in hospitals, and potential risks to patient safety. The full list of the symptoms and the exact prompts used is given in the Appendix N.
@@ 5 Prescriptive component in concept prototypes
One of the basic characteristics of System-1 is that it represents concepts with prototypical examples Kahneman (2011). In humans, though a prototype is often understood as the most typical/representative member of a concept (Murphy, 2004), they are found to embody both statistical regularities and goal-oriented ideals within a concept (Barsalou, 1985).
For instance, a ‘Robin’ might be considered a prototype of the concept ‘Bird’, as it shares many common features with most birds with high occurrence (statistics), and has the ability to fly (a value expected of birds), making it a prototypical example of the concept ‘bird’ (Smith and Medin, 1981). For this reason, penguin-a flightless bird, is less prototypical bird than ‘Robin’. Prototypicality defines normality of a concept that drive the sampling (Appendix  F).
Unlike humans, it is not clear whether LLMs rely on concept prototypes for sampling. But since the sampling heuristics of LLMs converge with humans, it is interesting to investigate concept prototypicality in LLMs. We do not claim that LLMs output is prototype driven, but make an initial exploration in this direction using the exact setting as in Bear and Knobe (2017).
We use eight concepts, and for each concept $C$, six different exemplars. Exemplars are short descriptions of items of a concept. For instance, for the concept of ‘High-school teacher’, the first exemplar is as follows: ‘A 30-year-old woman who basically knows the material she is teaching but is relatively uninspiring, boring to listen to, and not particularly fond of her job’.
Similar to experiment protocol in Bear and Knobe (2017), LLMs rate each exemplars on three dimensions: average, ideal, and the prototypicality of the exemplar. Prototypicality score is derived by averaging three entities, which measure the degree to which the given prototype is a “good example”, “paradigmatic example”, or “prototypical example” Bear and Knobe (2017). The LLM is asked to rate on a 7-point scale ranging from not at all average/ideal/‘good example’, which has a score of 0, to completely average/ideal/‘good example’, with a score 7. The full set of concepts and exemplars are in Appendix P.
@@
Evaluating across different LLMs, we obtain the following results: Llama-3-7b (binomial $p$ = 0.003), Mixtral-8x7B (binomial $p$ = 0.05), GPT3.5-turbo (binomial $p$ < 0.001), Claude (binomial $p$ < 0.001), Mistral (binomial $p$ = 0.0019), indicating the effect of prescriptive norms in prototypes of concepts. The complete set of results for every exemplar is given in Appendix Q. This experiment is an initial exploration, finding that LLMs’ concept of prototypes is influenced not only by statistical averages but also by an underlying prescriptive norm. These findings suggest that the LLM’s judgment of what constitutes a typical or prototypical example is systematically biased toward idealized representations calling for further investigations in this direction.